# An algorithmic version of the Hajnal–Szemerédi theorem

### v2

@@ 1.3. Sharpness of the result
We build the first parameterized algorithm for Problem 1.2 that bridges the gap from $c=0$ (P) to $c=n^{\Theta(1)}$ (NP-C). However, the running time $n^{O(1)}2^{O(c^{{r}^{r+4}})}$ of our parameterized algorithm can probably be improved – it will be interesting to design more efficient algorithms with reduced exponent of $c$, or even of type $n^{O(1)}2^{O(c)}$, which would yield polynomial-time solution for $c=O(\log n)$.
@@
For positive integers $a,b$ with $a<b$, let $[a]=\{1,2,\ldots,a\}$ and $[a,b]=\{a,a+1,\ldots,b\}$. For constants $x,y,z$, $x=y\pm z$ means that $y-z\leq x\leq y+z$. We write $x\ll y\ll z$ to mean that we choose constants from right to left, that is, for any $z>0$, there exist functions $f$ and $g$ such that, whenever $y\leq f(z)$ and $x\leq g(y)$, the subsequent statement holds. Hierarchies of other lengths are defined analogously.
@@ 2. Non-extremal case
We first apply Lemma 2.5 on $G$ and obtain in time $O(n^{4r^{2}})$ a $\xi$-absorbing set $A$ consisting of at most $\xi n$ vertex-disjoint $r^{2}$-sets such that every $r$-set has at least $\xi^{2}n$ absorbers in $A$. Let $n^{\prime}=|V(G)\setminus A)|$. Since $|A|\leq\xi r^{2}n$ and $\xi\ll\alpha$, we have $n^{\prime}\geq(1-\xi r^{2})n$ and $\delta(G-A)\geq(1-1/r-2\alpha)n\geq(1-1/r-2\alpha)n^{\prime}$. For an $\frac{n^{\prime}}{r}$-set in $V(G)\setminus A$, it induces at least $\gamma n^{2}-\xi rn^{2}\geq\gamma n^{2}/2$ edges and so $G-A$ has no $\frac{\gamma}{2}$-sparse $\frac{n^{\prime}}{r}$-set. Then we apply Lemma 2.6 to $G-A$ with $2\alpha$ in place of $\alpha$ and $\gamma/2$ in place of $\gamma$, and obtain in time $O(n^{r+1})$ a $K_{r}$-tiling $\mathcal{K}$ covering all but at most $\tau n^{\prime}$ vertices of $G-A$. Denote by $U$ the set of uncovered vertices in $G-A$. We arbitrarily divide $U$ into at most $\tau n^{\prime}/r$ vertex-disjoint $r$-sets. Since $\tau\ll\xi^{2}$, we can greedily assign each $r$-set a unique absorber in $A$. Hence $G[A\cup U]$ contains a $K_{r}$-factor. Together with $\mathcal{K}$, we obtain a $K_{r}$-factor of $G$. ∎
@@ 3. Absorbing lemma
[12, Proposition 4.7] Fix constants $\beta_{1}>\beta_{2}>0$ and integers $m,M,N$ and $\ell\leq N$ such that $N$ and $\ell$ are sufficiently large, and that $M\leq(1/8)\exp(\beta_{2}^{2}\ell/(3\beta_{1}))$. Let $U$ and $W$ be disjoint sets of sizes $|U|=M$ and $|W|=N$. Let $H$ be a graph with vertex set $U\cup W$ such that $H[U]$ is empty, $H[W]$ has exactly $m$ edges, and $d_{H}(u)\geq\beta_{1}N$ for every $u\in U$. Then we can find an independent set $L\subseteq W$ in $H$ in time $O(N^{4}+MN^{3})$ such that $(1-\nu)\ell\leq|L|\leq\ell$ and $|N_{H}(u)\cap L|\geq(\beta_{1}-\beta_{2}-\nu)\ell$ for every $u\in U$, where $\nu=2m\ell/N^{2}$.
@@
The proof follows the standard proof of the absorbing method, starting with the enumeration of the absorbers. We claim that any $r$-set $\{v_{1},\ldots,v_{r}\}\subset V(G)$ has $\Omega(n^{r^{2}})$ absorbers. Indeed, we first choose $r$ distinct vertices $u_{1},\ldots,u_{r}\in V(G)\setminus\{v_{1},\ldots,v_{r}\}$ one by one such that $\{u_{1},\ldots,u_{r}\}$ induces a copy of $K_{r}$. Since any $r-1$ vertices have at least $(r-1)(1-1/r-\alpha)n-(r-2)n\geq(1/r-\alpha r)n$ common neighbors in $G$, there are at least $(1/r-\alpha r)^{r}n^{r}/r!$ choices for $u_{1},\ldots,u_{r}$. Then for each $i\in[r]$ we choose $r-3$ distinct vertices $w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ from the common neighborhood of $v_{i}$ and $u_{i}$ such that both $v_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ and $u_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ induce copies of $K_{r-2}$. Finally, we choose an edge $x_{i}y_{i}$ from the common neighborhood $U_{i}$ of $v_{i},u_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ for each $i\in[r]$. Since $|U_{i}|\geq(1/r-\alpha r)n$ and $G$ has no $\gamma$-sparse $\frac{n}{r}$-set, $G[U_{i}]$ induces at least $(\gamma-\alpha r)n^{2}$ edges for any $i\in[r]$. Hence $\bigcup_{i=1}^{r}\{u_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}},x_{i},y_{i}\}$ is an absorber for the $r$-set $\{v_{1},\ldots,v_{r}\}$ and it has at least
@@ 4.1. Regularity
[20, Slicing lemma] Let $(X,Y)$ be an $\varepsilon$-regular pair with density $d$, and for some $\eta>\varepsilon$, let $X^{\prime}\subseteq X,Y^{\prime}\subseteq Y$ with $|X^{\prime}|\geq\eta|X|,|Y^{\prime}|\geq\eta|Y|$. Then $(X^{\prime},Y^{\prime})$ is an $\varepsilon^{\prime}$-regular pair with $\varepsilon^{\prime}=\max\{\varepsilon/\eta,2\varepsilon\}$, and for its density $d^{\prime}$ we have $|d^{\prime}-d|<\varepsilon$.
@@
[20, Counting lemma] Given $d>\varepsilon>0$, $m\in\mathbb{N}$ and $H$ some fixed graph on $r$ vertices, let $G$ be a graph obtained by replacing every vertex $x_{i}$ of $H$ with an independent set $V_{i}$ of size $m$ and every edge of $H$ with an $(\varepsilon,d)$-regular pair on the corresponding sets. If $\varepsilon\leq\frac{d^{r}}{(2+r)2^{r}}=:d_{0}$, then there are at least $(d_{0}m)^{r}$ embeddings of $H$ in $G$ so that each $x_{i}$ is embedded into the set $V_{i}$.
@@
[1, 20, Degree form of the Regularity Lemma] For every $\varepsilon>0$, there is an $M=M(\varepsilon)$ such that if $G=(V,E)$ is an $n$-vertex graph and $d\in[0,1]$ is any real number, then we can find in time $O(n^{2.376})$ a partition $V=V_{0}\cup V_{1}\cup\cdots\cup V_{k}$ and a spanning subgraph $G^{\prime}\subseteq G$ with the following properties:
@@
  • all pairs $(V_{i},V_{j})$ $(1\leq i<j\leq k)$ are $\varepsilon$-regular in $G^{\prime}$ with density 00 or at least $d$, where in the latter case, we indeed have $E_{G^{\prime}}(V_{i},V_{j})=E_{G}(V_{i},V_{j})$.
@@
Let $10<\beta<$ 1 be a real number. If $R$ has a $\beta$-sparse $\frac{k}{r}$-set, then $G$ has a $\left(\beta+2\varepsilon+d\right)$-sparse $\frac{n}{r}$-set.
@@
Let $10<\eta<$ 1. If $R$ has no $\left(\frac{8\eta}{\varepsilon^{2}d}+\frac{3\varepsilon}{r}\right)$-sparse $\frac{k}{r}$-set, then $G$ has no $\eta$-sparse $\frac{n}{r}$-set.
@@ 4.2. Almost perfect matching in $r$-complexes
In the proof of Lemma 2.6, we will use the result in [17] about the existence of almost perfect matching in an $r$-complex. To state the results, we first introduce some definitions. An $r$-system is a hypergraph $J$ in which every edge of $J$ has size at most $r$ and $\emptyset\in J$. An edge of size $k$ ($k\leq r$) in $J$ is called a $k$-edge of $J$. We use $J_{k}$ to denote the $k$-uniform hypergraph on $V(J)$ with all $k$-edges of $J$.
@@
Suppose $V$ is a set of $n$ vertices, $j\in[r-1]$ and $S\subseteq V$. Let $J=J(S,j)$ be the $r$-complex in which $J_{i}$ (for $0\leq i\leq r$) consists of all $i$-sets in $V$ that contains at most $j$ vertices of $S$.
@@
Suppose $G$ and $H$ are $r$-uniform graphs on the same set $V$ of $n$ vertices and $10<\beta<$ 1. We say that $G$ is $\beta$-contained in $H$ if all but at most $\beta n^{k}$ edges of $G$ are edges of $H$.
@@
Proof.
Given $\gamma_{H}\in(0,1)$, we choose $\delta,\varepsilon\ll\varepsilon_{1}\ll\varepsilon_{2}\ll\gamma_{H}$. Suppose that $H$ has less than $\delta h^{k}$ copies of $K_{k}$. By the removal lemma [3], $H$ can be made $K_{k}$-free by removing at most $\varepsilon_{1}h^{2}$ edges. By Erdős–Simonovits stability theorem, the remaining graph can be changed to the Turán graph $T_{k-1}(h)$ by altering at most $\varepsilon_{2}h^{2}$ adjacencies. Hence $H$ has a $\gamma_{H}$-sparse $\frac{h}{k-1}$-set, a contradiction. ∎
We give the following corollary, which is easier to apply than Theorem 4.12.
@@
(${r\choose{j+1}}\beta n^{r}\bigg{/}\left(\frac{n}{2r^{2}}\right)^{r-j-1}\leq(2r^{2})^{r}\beta n^{j+1}$
@@ 5. Preparation for the extremal case
The following definitions are key to our proofs, and motivated by earlier works on problems of similar flavor. The idea behind is that when $G$ contains large sparse sets, one can find and analyze its (rough) structure.
All vertex partitions considered in this paper have an implicit order on their parts.
Definition 5.1 (Index vector & Slack).
@@
Let $r\geq r^{\prime}\geq 2$ and $\varepsilon<1/r$. Let $t_{1},t_{2},\ldots,t_{r^{\prime}}$ be $r^{\prime}$ integers such that $|t_{i}|\leq\varepsilon n$ for each $i\in[r^{\prime}]$ and $\sum_{i=1}^{r^{\prime}}t_{i}=0$. Let $G$ be an $n$-vertex graph and $(V_{1},V_{2},\dots,V_{r^{\prime}})$ be a partition of $V(G)$ with $|V_{i}|=n/r+t_{i}$ for each $i\in[r^{\prime}-1]$ and $|V_{r^{\prime}}|=(r-r^{\prime}+1)n/r+t_{r^{\prime}}$. Let $\textbf{b}=(1,\ldots,1,r-r^{\prime}+1)$ be an $r^{\prime}$-dimensional vector. For any $i\in[r^{\prime}]$, let $\textbf{u}_{i}$ be the $i$-th unit vector, i.e. $\textbf{u}_{i}$ has 1 on the $i$-th coordinate and 0 on the other coordinates. Let $F$ be a $K_{r}$-tiling of size $f$ in $G$.
@@
  • The index vector of $F$ is $\textbf{i}(F)=(x_{1},x_{2},\ldots,x_{r^{\prime}})$ by choosing $x_{i}=|V(F)\cap V_{i}|$ for every $i\in[r^{\prime}]$.
  • The slack of $F$ is $\textbf{s}(F)=(s_{1},s_{2},\dots,s_{r^{\prime}}):=\textbf{i}(F)-f\textbf{b}$. Note that $\sum_{i=1}^{r^{\prime}}s_{i}=0$ and $\sum_{s_{i}>0}s_{i}\leq(r-1)f$.
  • The slack of the partition $(V_{1},V_{2},\dots,V_{r^{\prime}})$ is $(t_{1},t_{2},\dots,t_{r^{\prime}})=(|V_{1}|,|V_{2}|,\dots,|V_{r^{\prime}}|)-\frac{n}{r}\textbf{b}$. The partition $(V_{1},V_{2},\dots,V_{r^{\prime}})$ is balanced if $|V_{1}|=\cdots=|V_{r^{\prime}-1}|=|V_{r^{\prime}}|/{(r-r^{\prime}+1)}$.
  • A copy of $K_{r}$ in $G$ is balanced if its index vector is b. Compared with a balanced copy of $K_{r}$, an $(i,j)$-copy of $K_{r}$ has index vector $\textbf{b}+\textbf{u}_{i}-\textbf{u}_{j}$, where $i,j\in[r^{\prime}]$ are distinct.
@@
  • For a copy of $K_{r}$, we refer to its index vector as its type. Note that there are $y:={{r+r^{\prime}-1}\choose{r^{\prime}-1}}$ different types of $K_{r}$.
  • Let $I$ be a subset of $[r^{\prime}]$ and v be an $r^{\prime}$-dimensional vector. We use $\textbf{v}|_{I}$ to denote the $|I|$-dimensional vector obtained by restricting v to the coordinates in $I$.
@@
The definitions are natural descriptions of the vertex distributions of a $K_{r}$-tiling. The key idea here is that the slack of a $K_{r}$-factor (which covers $V(G)=V_{1}\cup V_{2}\cup\cdots\cup V_{r^{\prime}}$) is equal to the slack of the partition, each coordinate of which is small. Then, a necessary condition for having a $K_{r}$-factor is to have a $K_{r}$-tiling with the same slack. The question is how small we can take such a $K_{r}$-tiling, which is answered by Lemma 1.7. Below we restate it here (using slack) for convenience and give its proof.
@@
Lemma 1.7. Let $r\geq r^{\prime}\geq 2$, $I\subseteq[r^{\prime}]$. Let $t_{1},t_{2},\ldots,t_{r^{\prime}}$ be $r^{\prime}$ integers such that $\sum_{i=1}^{r^{\prime}}t_{i}=0$ and $\sum_{i\in I}|t_{i}|>4r-3$. Let $G$ be an $n$-vertex graph and $(V_{1},V_{2},\dots,V_{r^{\prime}})$ be a partition of $V(G)$ with $|V_{i}|=n/r+t_{i}$ for each $i\in[r^{\prime}-1]$ and $|V_{r^{\prime}}|=(r-r^{\prime}+1)n/r+t_{r^{\prime}}$. Suppose that $G$ contains a $K_{r}$-factor $K$ whose slack is $\textbf{s}(K)=(t_{1},t_{2},\dots,t_{r^{\prime}})$. Denote $t_{I}=\sum_{i\in I}|t_{i}|$. Then we can find a subset $K^{\prime}\subseteq K$ of size at most $(3t_{I})^{r}$ such that $\textbf{s}(K^{\prime})|_{I}=\textbf{s}(K)|_{I}$. In particular, $G-K^{\prime}$ still contains a $K_{r}$-factor.

### v3

@@ 1.3. Sharpness of the result
We build the first parameterized algorithm for $K_{r}$-Factor In Graphs With High Minimum Degree that bridges the gap from $c=0$ (P) to $c=n^{\Theta(1)}$ (NP-C). However, the running time $n^{O(1)}2^{O(c^{{r}^{r+4}})}$ of our parameterized algorithm can probably be improved – it will be interesting to design more efficient algorithms with reduced exponent of $c$, or even of type $n^{O(1)}2^{O(c)}$ (we have already achieved this type for $r=3$), which would yield a polynomial-time solution for $c=O(\log n)$.
@@
For positive integers $a,b$ with $a<b$, let $a]=\{1,2,\ldots,a\}$ and $a,b]=\{a,a+1,\ldots,b\}$. For constants $x,y,z$, $x=y\pm z$ means that $y-z\leq x\leq y+z$. We write $x\ll y\ll z$ to mean that we choose constants from right to left, that is, for any $z>0$, there exist functions $f$ and $g$ such that, whenever $y\leq f(z)$ and $x\leq g(y)$, the subsequent statement holds. Hierarchies of other lengths are defined analogously.
@@ 2. Non-extremal case
We first apply Lemma 2.5 on $G$ and obtain in time $O(n^{4r^{2}})$ a $\xi$-absorbing set $A$ consisting of at most $\xi n$ vertex-disjoint $r^{2}$-sets such that every $r$-set has at least $\xi^{2}n$ absorbers in $A$. Let $n^{\prime}=|V(G)\setminus A|$. Since $|A|\leq\xi r^{2}n$ and $\xi\ll\alpha$, we have $n^{\prime}\geq(1-\xi r^{2})n$ and $\delta(G-A)\geq(1-1/r-2\alpha)n\geq(1-1/r-2\alpha)n^{\prime}$. For an $\frac{n^{\prime}}{r}$-set in $V(G)\setminus A$, it induces at least $\gamma n^{2}-\xi rn^{2}\geq\gamma n^{2}/2$ edges and so GAG-A has no $\frac{\gamma}{2}$-sparse $\frac{n^{\prime}}{r}$-set. Then we apply Lemma 2.6 to GAG-A with $2\alpha$ in place of $\alpha$ and $\gamma/2$ in place of $\gamma$, and obtain in time $O(n^{r+1})$ a $K_{r}$-tiling $\mathcal{K}$ covering all but at most $\tau n^{\prime}$ vertices of GAG-A. Denote by $U$ the set of uncovered vertices in GAG-A. We arbitrarily divide $U$ into at most $\tau n^{\prime}/r$ vertex-disjoint $r$-sets. Since $\tau\ll\xi^{2}$, we can greedily assign each $r$-set a unique absorber in $A$. Hence $G[A\cup U]$ contains a $K_{r}$-factor. Together with $\mathcal{K}$, we obtain a $K_{r}$-factor of $G$. ∎
@@ 3. Absorbing lemma
[15, Proposition 4.7] Fix constants $\beta_{1}>\beta_{2}>0$ and integers $m,M,N$ and $\ell\leq N$ such that $N$ and $\ell$ are sufficiently large, and that $M\leq(1/8)\exp(\beta_{2}^{2}\ell/(3\beta_{1}))$. Let $U$ and $W$ be disjoint sets of sizes $|U|=M$ and $|W|=N$. Let $H$ be a graph with vertex set $U\cup W$ such that $H[U]$ is empty, $H[W]$ has exactly $m$ edges, and $d_{H}(u)\geq\beta_{1}N$ for every $u\in U$. Then we can find an independent set $L\subseteq W$ in $H$ in time $O(N^{4}+MN^{3})$ such that $(1-\nu)\ell\leq|L|\leq\ell$ and $|N_{H}(u)\cap L|\geq(\beta_{1}-\beta_{2}-\nu)\ell$ for every $u\in U$, where $\nu=2m\ell/N^{2}$.
@@
The proof follows the standard proof of the absorbing method, starting with the enumeration of the absorbers. We claim that any $r$-set $\{v_{1},\ldots,v_{r}\}\subset V(G)$ has $\Omega(n^{r^{2}})$ absorbers. Indeed, we first choose $r$ distinct vertices $u_{1},\ldots,u_{r}\in V(G)\setminus\{v_{1},\ldots,v_{r}\}$ one by one such that $\{u_{1},\ldots,u_{r}\}$ induces a copy of $K_{r}$. Since any r1r-1 vertices have at least $(r-1)(1-1/r-\alpha)n-(r-2)n\geq(1/r-\alpha r)n$ common neighbors in $G$, there are at least $(1/r-\alpha r)^{r}n^{r}/r!$ choices for $u_{1},\ldots,u_{r}$. Then for each $i\in[r]$ we choose r3r-3 distinct vertices $w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ from the common neighborhood of $v_{i}$ and $u_{i}$ such that both $v_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ and $u_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ induce copies of $K_{r-2}$. Finally, we choose an edge $x_{i}y_{i}$ from the common neighborhood $U_{i}$ of $v_{i},u_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}}$ for each $i\in[r]$. Since $|U_{i}|\geq(1/r-\alpha r)n$ and $G$ has no $\gamma$-sparse $\frac{n}{r}$-set, $G[U_{i}]$ induces at least $(\gamma-\alpha r)n^{2}$ edges for any $i\in[r]$. Hence $\bigcup_{i=1}^{r}\{u_{i},w_{i1},w_{i2},\ldots,w_{i,{r-3}},x_{i},y_{i}\}$ is an absorber for the $r$-set $\{v_{1},\ldots,v_{r}\}$ and it has at least
@@ 4.1. Regularity
[24, Slicing lemma] Let $(X,Y)$ be an $\varepsilon$-regular pair with density $d$, and for some $\eta>\varepsilon$, let $X^{\prime}\subseteq X,Y^{\prime}\subseteq Y$ with $|X^{\prime}|\geq\eta|X|,|Y^{\prime}|\geq\eta|Y|$. Then $(X^{\prime},Y^{\prime})$ is an $\varepsilon^{\prime}$-regular pair with $\varepsilon^{\prime}=\max\{\varepsilon/\eta,2\varepsilon\}$, and for its density $d^{\prime}$ we have $|d^{\prime}-d|<\varepsilon$.
@@
[24, Counting lemma] Given $d>\varepsilon>0$, $m\in\mathbb{N}$ and $H$ some fixed graph on $r$ vertices, let $G$ be a graph obtained by replacing every vertex $x_{i}$ of $H$ with an independent set $V_{i}$ of size $m$ and every edge of $H$ with an $(\varepsilon,d)$-regular pair on the corresponding sets. If $\varepsilon\leq\frac{d^{r}}{(2+r)2^{r}}=:d_{0}$, then there are at least $(d_{0}m)^{r}$ embeddings of $H$ in $G$ so that each $x_{i}$ is embedded into the set $V_{i}$.
@@
[1, 24, Degree form of the Regularity Lemma] For every $\varepsilon>0$, there is an $M=M(\varepsilon)$ such that if $G=(V,E)$ is an $n$-vertex graph and $d\in[0,1]$ is any real number, then we can find in time $O(n^{2.376})$ a partition $V=V_{0}\cup V_{1}\cup\cdots\cup V_{k}$ and a spanning subgraph $G^{\prime}\subseteq G$ with the following properties:
@@
  • all pairs $(V_{i},V_{j})$ $(1\leq i<j\leq k)$ are $\varepsilon$-regular in $G^{\prime}$ with density 0 or at least $d$, where in the latter case, we indeed have $E_{G^{\prime}}(V_{i},V_{j})=E_{G}(V_{i},V_{j})$.
@@
Let $0<\beta<1$ be a real number. If $R$ has a $\beta$-sparse $\frac{k}{r}$-set, then $G$ has a $\left(\beta+2\varepsilon+d\right)$-sparse $\frac{n}{r}$-set.
@@
Let $0<\eta<1$. If $R$ has no $\left(\frac{8\eta}{\varepsilon^{2}d}+\frac{3\varepsilon}{r}\right)$-sparse $\frac{k}{r}$-set, then $G$ has no $\eta$-sparse $\frac{n}{r}$-set.
@@ 4.2. Almost perfect matching in $r$-complexes
In the proof of Lemma 2.6, we will use the result in [21] about the existence of an almost perfect matching in an $r$-complex. To state the results, we first introduce some definitions. An $r$-system is a hypergraph $J$ in which every edge of $J$ has size at most $r$ and $\emptyset\in J$. An edge of size $k$ ($k\leq r$) in $J$ is called a $k$-edge of $J$. We use $J_{k}$ to denote the $k$-uniform hypergraph on $V(J)$ with all $k$-edges of $J$.
@@
Suppose $V$ is a set of $n$ vertices, $j\in[r-1]$ and $S\subseteq V$. Let $J(S,j)$ be the $r$-complex in which $J(S,j)_{i}$ (for $0\leq i\leq r$) consists of all $i$-sets in $V$ that contain at most $j$ vertices of $S$.
@@
Suppose $G$ and $H$ are $r$-uniform graphs on the same set $V$ of $n$ vertices and $0<\beta<1$. We say that $G$ is $\beta$-contained in $H$ if all but at most $\beta n^{r}$ edges of $G$ are edges of $H$.
@@
We will give a corollary of Theorem 4.12, i.e. Corollary 4.15, which is easier to apply than Theorem 4.12. The proof of Corollary 4.15 also use the following fact.
@@
To prove Fact 4.13, we shall use the graph removal lemma by Erdős, Frankl and Rödl [7] and the Erdős–Simonovits stability theorem [6].
Lemma 4.14.
[7] For any graph $F$ and any $\epsilon>0$, there exists $\alpha>0$ such that any graph on $n$ vertices which contains at most $\alpha n^{v(F)}$ copies of $F$ may be made $F$-free by removing at most $\epsilon n^{2}$ edges.
A classical result in extremal graph theory is the Erdős–Simonovits stability theorem [6], which says that for every $\epsilon>0$ there exists $\alpha>0$ such that if $G$ is a $K_{r+1}$-free graph on $n$ vertices with $e(G)\geq\mathrm{ex}(n,K_{r+1})-\alpha n^{2}$, then $G$ can be obtained from the Turán graph $T_{k}(n)$ by adding and deleting at most $\epsilon n^{2}$ edges. Now we are ready to prove Fact 4.13.
Proof of Fact 4.13.
Given $\gamma_{H}\in(0,1)$, we choose $\delta,\varepsilon\ll\varepsilon_{1}\ll\varepsilon_{2}\ll\gamma_{H}$. Suppose that $H$ has less than $\delta h^{k}$ copies of $K_{k}$, by Lemma 4.14 with $\epsilon=\varepsilon_{1}$ and $\alpha=\delta$, $H$ can be made $K_{k}$-free by removing at most $\varepsilon_{1}h^{2}$ edges. By Erdős–Simonovits stability theorem [6] with $\epsilon=\varepsilon_{2}$ and $\alpha=\varepsilon/2$, the remaining graph can be changed to the Turán graph $T_{k-1}(h)$ by altering at most $\varepsilon_{2}h^{2}$ adjacencies. Hence $H$ has a $\gamma_{H}$-sparse $\frac{h}{k-1}$-set, a contradiction. ∎
Now we state the corollary of Theorem 4.12 and give its proof.
@@
${r\choose{j+1}}\beta n^{r}\bigg/\left(\frac{n}{2r^{2}}\right)^{r-j-1}\leq(2r^{2})^{r}\beta n^{j+1}$
@@ 5. Preparation for the extremal case
The following definitions are key to our proofs, and motivated by earlier works on problems of similar flavor. The idea behind it is that when $G$ contains large sparse sets, one can find and analyze its (rough) structure. All vertex partitions considered in this paper have an implicit order on their parts.
@@
Let $r\geq r^{\prime}\geq 2$ and $\varepsilon<1/r$. Let $t_{1},t_{2},\ldots,t_{r^{\prime}}$ be $r^{\prime}$ integers such that $|t_{i}|\leq\varepsilon n$ for each $i\in[r^{\prime}]$ and $\sum_{i=1}^{r^{\prime}}t_{i}=0$. Let $G$ be an $n$-vertex graph and $(V_{1},V_{2},\dots,V_{r^{\prime}})$ be a partition of $V(G)$ with $|V_{i}|=n/r+t_{i}$ for each $i\in[r^{\prime}-1]$ and $|V_{r^{\prime}}|=(r-r^{\prime}+1)n/r+t_{r^{\prime}}$. Let $\boldsymbol{b}=(1,\ldots,1,r-r^{\prime}+1)$ be an $r^{\prime}$-dimensional vector. For any $i\in[r^{\prime}]$, let $\boldsymbol{u}_{i}$ be the $i$-th unit vector, i.e. $\boldsymbol{u}_{i}$ has 1 on the $i$-th coordinate and 0 on the other coordinates. Let $F$ be a $K_{r}$-tiling of size $f$ in $G$.
Definition 5.1 (Index vector).
@@
The index vector of $F$ is $\textbf{i}(F)=(x_{1},x_{2},\ldots,x_{r^{\prime}})$ by choosing $x_{i}=|V(F)\cap V_{i}|$ for every $i\in[r^{\prime}]$. A copy of $K_{r}$ in $G$ is balanced if its index vector is b.
@@
Compared with a balanced copy of $K_{r}$, an $(i,j)$-copy of $K_{r}$ has index vector $\boldsymbol{b}+\boldsymbol{u}_{i}-\boldsymbol{u}_{j}$, where $i,j\in[r^{\prime}]$ are distinct. For a copy of $K_{r}$, we refer to its index vector as its type. Note that there are $y:={{r+r^{\prime}-1}\choose{r^{\prime}-1}}$ different types of $K_{r}$.
Definition 5.2 (Slack).
The slack of $F$ is $\textbf{s}(F)=(s_{1},s_{2},\dots,s_{r^{\prime}}):=\textbf{i}(F)-f\textbf{b}$. Note that $\sum_{i=1}^{r^{\prime}}s_{i}=0$ and $\sum_{s_{i}>0}s_{i}\leq(r-1)f$. The slack of the partition $(V_{1},V_{2},\dots,V_{r^{\prime}})$ is $(t_{1},t_{2},\dots,t_{r^{\prime}})=(|V_{1}|,|V_{2}|,\dots,|V_{r^{\prime}}|)-\frac{n}{r}\textbf{b}$. The partition $(V_{1},V_{2},\dots,V_{r^{\prime}})$ is balanced if $|V_{1}|=\cdots=|V_{r^{\prime}-1}|=|V_{r^{\prime}}|/{(r-r^{\prime}+1)}$.
@@
The definitions are natural descriptions of the vertex distributions of a $K_{r}$-tiling. The key idea here is that the slack of a $K_{r}$-factor (which covers $V(G)=V_{1}\cup V_{2}\cup\cdots\cup V_{r^{\prime}}$) is equal to the slack of the partition, each coordinate of which is small. Then, a necessary condition for having a $K_{r}$-factor is to have a $K_{r}$-tiling with the same slack. The question is how small we can take such a $K_{r}$-tiling, which is answered by Lemma 1.8. Below we restate it here (using slack) for convenience and give its proof. Before proceeding, we introduce a piece of notation. Let $I$ be a subset of $r^{\prime}]$ and $\boldsymbol{v}$ be an $r^{\prime}$-dimensional vector. We use $\boldsymbol{v}|_{I}$ to denote the $|I|$-dimensional vector obtained by restricting $\boldsymbol{v}$ to the coordinates in $I$.
@@
Lemma 1.8. Let $r\geq r^{\prime}\geq 2$, $I\subseteq[r^{\prime}]$. Let $t_{1},t_{2},\ldots,t_{r^{\prime}}$ be $r^{\prime}$ integers such that $\sum_{i=1}^{r^{\prime}}t_{i}=0$ and $\sum_{i\in I}|t_{i}|>4r-3$. Let $G$ be an $n$-vertex graph and $(V_{1},V_{2},\dots,V_{r^{\prime}})$ be a partition of $V(G)$ with $|V_{i}|=n/r+t_{i}$ for each $i\in[r^{\prime}-1]$ and $|V_{r^{\prime}}|=(r-r^{\prime}+1)n/r+t_{r^{\prime}}$. Suppose that $G$ contains a $K_{r}$-factor $K$ whose slack is $\textbf{s}(K)=(t_{1},t_{2},\dots,t_{r^{\prime}})$. Denote $t_{I}=\sum_{i\in I}|t_{i}|$. Then we can find a subset $K^{\prime}\subseteq K$ of size at most $g(t_{I})$ such that $\textbf{s}(K^{\prime})|_{I}=\textbf{s}(K)|_{I}$, where $g(t_{I})=(3t_{I})^{r}$ for $r\geq 4$ and $g(t_{I})=8t_{I}$ for $r=3$. In particular, $G-K^{\prime}$ still contains a $K_{r}$-factor.
Note that in this lemma we require the conclusion to hold for any given $I$, and that the $K^{\prime}$ only adjusts the slack of the parts corresponding to $I$. Indeed, $K^{\prime}$ will be searched by the color-coding technique (Lemma 5.4). If this lemma only holds for $I=[r^{\prime}]$, then the slack of $K^{\prime}$ equals to the slack of the whole partition. However, the running time required to search for $K^{\prime}$ might become larger than expected, that is, no longer being polynomial. In the proof of the extremal case, we first classify the parts by the coordinates of the slack, and let $I\subseteq[r^{\prime}]$ be the set of indices with small (to be quantified) slack coordinates. In order to find a $K_{r}$-tiling $H$ whose slack is equal to the slack of the partition, we first use color-coding to find a $K_{r}$-tiling $K^{\prime}$ that adjusts the slack of the parts corresponding to $I$. If no such structure can be found, then $G$ contains no $K_{r}$-factor by the contrapositive of Lemma 1.8. Otherwise, we will proceed to construct the remaining part of $H$. This remaining part is guaranteed to exist and can be found in polynomial time.